OpenStudy (anonymous):

James has twelve coins in his piggy bank. Some of the coins are quarters, some are nickels, and he has a total of $3.25. Which system of equations can be used to determine how many quarters, x, and how many nickels, y, he has?

OpenStudy (texaschic101):

x + y = 12 .25x + .05y = 3.25 or x + y = 12 25x + 5y = 325

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

anytime :)